OpenStudy (vinzkie2109):

If b is the mean proportional between a and c such that a:b=b:c prove that: abc(a+b+c)^3 = (ab+bc+ac)^3

OpenStudy (supersmart1001):

hmm...

OpenStudy (mathstudent55):

I got it.

OpenStudy (mathstudent55):

\(a:b = b:c\) \(b^2 = ac\) \(b = \sqrt{ac} \) \(abc(a + b + c )^3 =\) \(=ac\sqrt{ac}(a + b + c)^3\) \(= \sqrt{ac}\sqrt{ac}\sqrt{ac}(a + b + c)^3\) \(= \sqrt{ac}\sqrt{ac}\sqrt{ac}(a + b + c)(a + b + c)(a + b + c)\) \(= (a\sqrt{ac} + b\sqrt{ac} + c\sqrt{ac})^3\) \(= (ab + b^2 + bc)^3\) \(= (ab + ac + bc)^3\) \(= (ab + bc + ac )^3\)

OpenStudy (mathstudent55):

@vinzkie2109 Do you understand it?

OpenStudy (vinzkie2109):

thanks a lot!